OpenStudy (anonymous):

Choose the correct equation in point-slope form for the line through the given point with the given slope. (4, –6) M=3/5

jigglypuff314 (jigglypuff314):

point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1)